OpenStudy (anonymous):

if f(x)=-2x-1 and g(x)=6x-1 find f[g(-20]

OpenStudy (anonymous):

f(g(x))=-2(6x-1)-1,f(g(-20))=-2(6*-20-1)=242

OpenStudy (anonymous):

g(-20)=6(-20)-1=-121 f(g(-20)=f(-121)=-2(-121)-1=241 @ASAAD123 you forgot the "-1"

OpenStudy (anonymous):

yup @Zantout